OpenStudy (anonymous):

The equation of a line is y=-2/3x-1 . What is the equation of the line that is perpendicular to the first line and passes through (–4, 3)?

OpenStudy (dmb1215):

ok give me a sec and i will have it for you

OpenStudy (texaschic101):

do you know what the slope of y = -2/3x - 1 is ?

OpenStudy (dmb1215):

its -2/3

OpenStudy (texaschic101):

correct. But we are looking for a perpendicular slope, so we need the negative reciprocal of that. All that means is " flip " the slope and change the sign. For example, the negative reciprocal of 3/4 is -4/3 (see how I flipped the slope and changed the sign). So what is the negative reciprocal of -2/3 ?

OpenStudy (dmb1215):

ok here is what i got 2x+3y=-3

OpenStudy (dmb1215):

so its y=-3/2x-3 wouldnt it

OpenStudy (texaschic101):

y = mx + b slope(m) = 3/2 (-4,3) x = -4 and y = 3 now we sub 3 = 3/2(-4) + b 3 = - 6 + b 3 + 6 = b 9 = b so your parallel equation is : y = 3/2x + 9 ==> slope intercept form but if you need it in standard form, it is : y = 3/2x + 9 -- multiply equation by 2 2y = 3x + 18 -- subtract 3x -3x + 2y = 18 -- multiply by -1 to make x positive 3x - 2y = -18 ==> standard form

OpenStudy (dmb1215):

so i had the letters in the wrong place

OpenStudy (texaschic101):

the slope you got was incorrect, as well.

OpenStudy (anonymous):

whats the answer guys ??

OpenStudy (texaschic101):

does anyone have any questions ? answer is : y = 3/2x + 9 in slope intercept form or 3x - 2y = -18, if you need it in standard form.

OpenStudy (anonymous):

i have alot of guestions can you answer them for me please ? and thx

OpenStudy (texaschic101):

ask away

OpenStudy (anonymous):

What is the equation of the line that passes through (–2, –3) and is perpendicular to 2x – 3y = 6?

OpenStudy (texaschic101):

same type of problem. Find the slope of the equation given by putting the equation in y = mx + b form, where m is your slope. 2x - 3y = 6 -- subtract 2x -3y = -2x + 6 -- divide by -3 y = 2/3x - 2 so our slope is 2/3. Our perpendicular slope will then be -3/2. y = mx + b slope(m) = -3/2 (-2,-3) x = -2 and y = -3 now sub -3 = -3/2(-2) + b -3 = 3 + b -3 - 3 = b -6 = b so your perpendicular equation is : y = -3/2x - 6

OpenStudy (anonymous):

another one he graph of which of the following will be perpendicular to the graph of y = 2x – 4? A. 2x + y = 4 B. 4x – 2y = 4 C. y = –2x + 8 D. 2x + 4y = 1

OpenStudy (texaschic101):

y = 2x - 4 --- slope is 2 we are looking for perpendicular, so we need slope of -1/2 I am going to put your answer choices in y = mx + b form, and I am looking for a slope of -1/2. (A) 2x + y = 4 y = -2x + 4 (incorrect) (B) 4x - 2y = 4 -2y = -4x + 4 y = 1/2x - 2 (incorrect) (C) y = -2x + 8 (incorrect) (D) 2x + 4y = 1 4y = -2x + 1 y = -1/2x + 1/4 (BINGO -- here it is) ANSWER D

OpenStudy (anonymous):

The vertices of a triangle are A(–2, 2), B(9, 0), and C(1, –4). If the triangle is a right triangle, which vertex is the right angle? A B C D. none, it is not a right triangle

OpenStudy (anonymous):

can you help me with that one ?

OpenStudy (texaschic101):

I am sorry...I do not know much about geometry :(